# An interpretation of the Brownian bridge as a physics-informed prior for the Poisson equation

### v1

@@ 6 Conclusions and outlook
In this work, we established a connection between the physics-informed machine learning approach for the Poisson equation and GP regression. Specifically, we showed the physics-informed loss function based on the variational form of the Poisson equation is a kernel method. Then, from the connections between kernel methods and GP regression, we showed that the loss function provides the MAP estimator for GP regression when starting with a Brownian bridge prior. In one-dimension, we may even move beyond GP regression and consider the prior as a Gaussian measure on $L^{2}(\Omega)$. This is in an effort to incorporate nonlinear measurement modalities into the framework.
Using the connection to GP regression, we studied different properties of the field reconstruction problem. In Section 4, we were able to prove convergence of the GP MAP estimator to the ground truth in the limit of infinite data. This also provides the result for the physics-informed optimization problem. We briefly discussed the consequences of this in the context of PINNs. We also derived a finite-dimensional basis representation of the prior as a subset of $L^{2}(\Omega)$. This is in contrast to the usual approach taken in GP regression, which instead learns the posterior on a mesh of $\Omega$. We proved in one-dimension that this representation converges to the correct Gaussian measure, and when $d>1$ we have convergence to the correct stochastic process in the setting of tempered distributions.
The main results of the paper are in Section 5, where we connect the method to the important problem of identifying model-form error. When we work under a physics-informed framework, we a priori assume the system is modeled by a specific form of the physics, which in this case is eq. (1.1). In any given application, it is entirely possible that we have picked the wrong model. The usual paradigm enforces the physics as a hard constraint and does not take this into account. We have modified the method so that the physics are enforced as a soft constraint. This is done through inclusion of the hyperparameter $\beta$.
@@
While the main focus of this work was on the Poisson equation, it is possible to extend the results to certain other PDEs. The main requirement is that the physics-informed loss function admits a quadratic form. This is so that it may be connected to a kernel method, from which we define a suitable GP. Another example one could study is the Helmholtz equation
$-\nabla^{2}u+\omega^{2}u+q=0.$ (6.1)
One can show that eq. (6.1) with Dirichlet boundaries has the energy functional [12]
@@
$E(u)=\frac{1}{2}\langle u-Cq,C^{-1}(u-Cq)\rangle,$
@@
Therefore, it appears this trick is limited to linear PDEs so that the Greens function may be identified. Also, in physics-informed machine learning, it is much more common to use the integrated square residual of the PDE to define the loss function, rather than a variational form. In a future work, we plan to extend this method both to nonlinear PDEs and to loss functions defined by the integrated square residual. This can be done via Taylor approximation.
Lastly, we restricted our work to theory, and did not touch on any numerical methods. While standard GP regression techniques may be used in applications, there are some computational issues which should be resolved. The main bottleneck is the fact that the mean function of the physics-informed prior is given by the solution to the PDE. If we are solving the inverse problem, then the mean function will change every time $q$ is updated, meaning that the PDE must be resolved. We plan to address this issue in future work by developing specialized sampling algorithms which avoid needing to call a PDE solver. This is based on the finite-dimensional basis representation derived in this work.
@@ Appendix A Gaussian measures
We summarize important concepts related to Gaussian measures on separable Hilbert spaces. Note that the theory of Gaussian measures on Hilbert spaces can easily be extended to the Banach space setting, but this is not needed in this work. The texts [32, 10] along with the notes provided in [20] provide a nice background to the theory.
@@
A linear operator $C:H\to H$ is said to be trace class if, for any orthonormal basis $\{\psi_{n}\}_{n=1}^{\infty}$ of $H$, we have
@@
[…]
@@
Theorem A.1 (Theorem 6.4 [44]).
@@
Theorem A.3 (Theorem 1 [35]).
@@
Appendix B Proof of Theorem 4.4
@@
Definition B.1 (Random field, Lévy white noise).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, and $U\subseteq\mathbb{R}^{d}$ an open set.
  1. (i)
    A random field $X$ on $U$ is a measurable mapping $X:U\times\Omega\to\mathbb{R}^{n}$ such that for any $x\in U$, $X(x;\cdot)$ is a real-valued random variable.
  2. (ii)
    A Lévy white noise $X$ (generalized random field) is a measurable mapping $u:(\Omega,\mathcal{F})\to\mathcal{S}^{\prime}(\mathbb{R}^{d})$. That is, $X(u)$ is a real-valued random variable $\forall u\in\mathcal{S}(\mathbb{R}^{d})$.
The characteristic function of a Lévy white noise is defined as
$\phi_{X}(u)=\mathbb{E}[\exp\{iX(u)\}]=\int_{\mathcal{S}^{\prime}(\mathbb{R}^{d})}\exp\{iL(u)\}\mu(dL),\quad u\in\mathcal{S}(\mathbb{R}^{d}).$
To apply Theorem 4.3, we must show that the finite-dimensional measure given by eq. (4.2) admits a Lévy white noise. To show this, we first must show the random field $X_{n}$ associated with each $\mu^{(n)}$ is linear in the following sense.
Definition B.2.
Let $X$ be a random field on $\mathcal{S}(\mathbb{R}^{d})$. We say that $X$ is linear if
$X\left(\sum_{i=1}^{m}\alpha_{i}u_{i}\right)=\sum_{i=1}^{m}\alpha_{i}X(u_{i}),\quad\mathrm{a.s.},$
for all $m\geq 1$, $\alpha_{1},\dots,\alpha_{m}\in\mathbb{R}$, and $u_{1},\dots,u_{m}\in\mathcal{S}(\mathbb{R}^{d})$.
The proof then relies on the following result.
Lemma B.1 (Corollary 2.2 [8]).
Let $X_{n}=(X_{n}(u))_{u\in\mathcal{S}(\mathbb{R}^{d})}$ be a collection of linear, real random variables on $(\Omega,\mathcal{F},\mathbb{P})$. If $\phi_{X_{n}}$ is continuous at 00, then there is a version of $X_{n}$ that is a Lévy white noise.
We can now prove
Lemma B.2.
For each $\mu^{(n)}$ as given by eq. (4.2), the associated random field $X_{n}$ admits a version which is a Lévy white noise.
@@
Restrict each $\mu^{(n)}$ to $\mathcal{S}(\mathbb{R}^{d})$, which can be done as $\mathcal{S}(\mathbb{R}^{d})\subset L^{2}(\mathbb{R}^{d})$. Then, each $\mu^{(n)}$ is a Gaussian measure associated with a Gaussian random field $X_{n}$ on $\mathcal{S}(\mathbb{R}^{d})$. Since $X_{n}$ is Gaussian, it has characteristic function
$\phi_{X_{n}}(u)=\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}}\hat{u}\rangle\right\},\quad u\in\mathcal{S}(\mathbb{R}^{d}),$
where $\hat{u}=\sum_{j=1}^{n}\langle u,\psi_{j}\rangle\psi_{j}$. It is obvious that $\phi_{X_{n}}$ is continuous at 00.
To show linearity, fix $m\geq 1$, $\alpha_{1},\dots,\alpha_{m}\in\mathbb{R}$, and $u_{1}\dots,u_{m}\in\mathcal{S}(\mathbb{R}^{d})$. We will show the random variables $X_{n}\left(\sum_{j=1}^{m}\alpha_{j}u_{j}\right)$ and $\sum_{j=1}^{m}\alpha_{j}X_{n}(u_{j})$ have the same characteristic functions and therefore share the same distribution. Now, the random variable $X_{n}\left(\sum_{j=1}^{m}\alpha_{j}u_{j}\right)$ is Gaussian, so it has characteristic function
$\phi_{X_{n}}\left(\sum_{j=1}^{m}\alpha_{j}u_{j}\right)$ $=\exp\left\{-\frac{1}{2}\left\langle\sum_{j=1}^{m}\alpha_{j}\hat{u}_{j},\Sigma_{\mathcal{F}}\left(\sum_{j=1}^{m}\alpha_{j}\hat{u}_{j}\right)\right\rangle\right\}$
$=\exp\left\{-\frac{1}{2}\sum_{j=1}^{m}\langle\alpha_{j}\hat{u}_{j},\Sigma_{\mathcal{F}}(\alpha_{j}\hat{u}_{j})\rangle\right\}$
$=\prod_{j=1}^{m}\exp\left\{-\frac{1}{2}\langle\alpha_{j}\hat{u}_{j},\Sigma_{\mathcal{F}}(\alpha_{j}\hat{u}_{j})\rangle\right\}$
$=\prod_{j=1}^{m}\phi_{X_{n}}(\alpha_{j}u_{j}).$
Recall the identity of characteristic functions $\alpha_{j}\phi_{X_{n}}(u_{j})=\phi_{X_{n}}(\alpha_{j}u_{j})$. So, $\sum_{j=1}^{m}\alpha_{j}X_{n}(u_{j})$ has characteristic function
$\mathbb{E}\left[\exp\left\{i\sum_{j=1}^{m}\alpha_{j}X_{n}(u_{j})\right\}\right]$ $=\mathbb{E}\left[\prod_{j=1}^{m}\exp\{i\alpha_{j}X_{n}(u_{j})\}\right]$
$=\prod_{j=1}^{m}\mathbb{E}[\exp\{i\alpha_{j}X_{n}(u_{j})\}]$
$=\prod_{j=1}^{m}\phi_{X_{n}}(\alpha_{j}u_{j}),$
which is the same characteristic function as before. Hence, the collection $(X_{n}(u))_{u\in\mathcal{S}(\mathbb{R}^{d})}$ is linear, and application of Lemma B.1 completes the proof. ∎
Finally, we will look to apply Theorem 4.3. By Lemma B.2, we may regard each $\mu^{(n)}$ in the sequence as a Lévy white noise. As they are Gaussian, each has characteristic function
$\phi_{X_{n}}(u)=\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}}\hat{u}\rangle\right\},\quad u\in\mathcal{S}(\mathbb{R}^{d}).$
Let $\phi(u)=\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}$, which is continuous at 00. Then in the limit, for any $u\in\mathcal{S}(\mathbb{R}^{d})$, we have
@@
By Lévy’s continuity theorem, there exists a Lévy white noise $\mu$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ such that $\mu^{(n)}\overset{\mathrm{d}}{\rightarrow}\mu$. Finally, $\phi$ is the form of a characteristic function of a Gaussian random field on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$.

### v2

@@ 6 Conclusions and outlook
In this work, we established a connection between the variational formulation of the Poisson equation and GP regression. Specifically, we showed the resulting inverse problem can be viewed as a kernel method. Then, from the connections between kernel methods and GP regression, we showed that the loss function provides the MAP estimator for GP regression when starting with a Brownian bridge prior. In one-dimension, we may even move beyond GP regression and consider the prior as a Gaussian measure on $L^{2}(\Omega)$, and on the dual of a Sobolev space in the general case. This is in an effort to incorporate nonlinear measurement modalities into the framework.
Using the connection to GP regression, we studied different properties of the field reconstruction problem. In Section 4, we were able to prove convergence of the GP MAP estimator to the ground truth in the limit of infinite data. This also provides the result for the related variational problem. We briefly discussed the consequences of this in the context of PINNs. We also derived a finite-dimensional basis representation of the prior as a subset of $L^{2}(\Omega)$. This is in contrast to the usual approach taken in GP regression, which instead learns the posterior on a mesh of $\Omega$. We proved that this representation and the corresponding posterior approximation converge to the correct Gaussian measure.
The main results of the paper are in Section 5, where we connect the method to the important problem of identifying model-form error. When we work under a physics-informed framework, we a priori assume the system is modeled by a specific form of the physics, which in this case is eq. (1.1). In any given application, it is entirely possible that we have picked the wrong model. The usual paradigm enforces the physics as a hard constraint and does not take this into account. We have modified the method so that the physics is enforced as a soft constraint. This is done through inclusion of the hyperparameter $\beta$.
@@
While the main focus of this work was on the Poisson equation, it is possible to extend the results to certain other PDEs. The main requirement is that the PDE can be cast as a variational problem which admits a quadratic positive-definite form. This is so that it may be connected to a kernel method, from which we define a suitable GP. Another example one could study is the Helmholtz equation
$\nabla^{2}u+\omega^{2}u+q=0.$ (6.1)
One can show that eq. (6.1) with Dirichlet boundary conditions has the energy functional [brezis2011functional]
@@
$E(u)=\frac{1}{2}\langle u-Cq,C^{-1}(u-Cq)\rangle+\mathrm{const.},$
@@
Therefore, it appears this trick is limited to linear PDEs so that the Greens function may be identified. Also, in in much of the literature, it is common to use the integrated square residual of the PDE to define the loss function, rather than a variational form. In a future work, we plan to extend this method both to nonlinear PDEs and to loss functions defined by the integrated square residual. This can be done via Taylor approximation.
Lastly, we restricted our work to theory, and did not touch on any numerical methods. While standard GP regression techniques may be used in applications, there are some computational issues which should be resolved. The main bottleneck is the fact that the mean function of the physics-informed prior is given by the solution to the PDE. If we are solving the inverse problem, then the mean function will change every time $q$ is updated, meaning that the PDE must be resolved. Note that this is also the case in classical Bayesian inverse problems [stuart2010inverse], where the physics is enforced in the likelihood. We plan to address this issue in future work by developing specialized sampling algorithms which avoid needing to call a PDE solver. This is based on the finite-dimensional basis representation derived in this work.
@@ Appendix A Gaussian measures
We summarize important concepts related to Gaussian measures on separable Hilbert spaces. Note that the theory of Gaussian measures on Hilbert spaces can easily be extended to the Banach space setting, but this is not needed in this work. The texts [kuo2006gaussian, bogachev1998gaussian] along with the notes provided in [eldredge2016analysis] provide a nice background to the theory.
@@
A linear operator $C:H\to H$ is said to be trace class if, for any orthonormal basis $(\psi_{n})_{n=1}^{\infty}$ of $H$, we have
@@
$\mu\left(\{v\in H:\langle u,v\rangle\leq a\}\right)=\int_{-\infty}^{a}\frac{1}{\sqrt{2\pi\sigma_{u}}}\exp\left\{-\frac{1}{2\sigma_{u}}(x-m_{u})^{2}\right\}dx.$
@@
Theorem A.1 (Theorem 6.4 [stuart2010inverse]).
@@
Theorem A.3 (Theorem 1 [minh2024infinite]).
@@
We will also make use of the Wasserstein distance as a metric between probability measures when studying convergence. Unlike other probabilistic metrics, the Wasserstein distance does not require absolute continuity between distributions, making it a bit more general. The definition is given in terms of the metric on $H$, which we will take to be the metric induced by the norm $\|\cdot\|_{H}$. Let $\mathcal{P}_{p}(H)$ denote the space of all Borel probability measures on $H$ with finite $p$-th moment. Then, the $p$-Wassertein distance is defined as
$W_{p}(\mu,\nu)\coloneqq\inf_{\gamma\in\Gamma(\mu,\nu)}\left(\int_{H\times H}\|u-v\|_{H}^{p}\gamma(dudv)\right)^{1/p},$
for two probability measures $\mu,\nu\in\mathcal{P}_{p}(H)$ and $\Gamma(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. In the case that the measures are Gaussian, there exists a useful identity, known as Gelbrich’s formula [gelbrich1990formula, Theorem 3.5]. This states that for $\mu=\mathcal{N}(m_{1},C_{1})$ and $\nu=\mathcal{N}(m_{2},C_{2})$ (both on $H$), the $2$-Wasserstein distance is given by
$W_{2}(\mu,\nu)=\sqrt{\|m_{1}-m_{2}\|^{2}_{H}+\mathrm{tr}\left(C_{1}\right)+\mathrm{tr}\left(C_{2}\right)-2\mathrm{tr}\left(\sqrt{C_{1}^{1/2}C_{2}C_{1}^{1/2}}\right)}.$ (A.1)
Appendix B Proof of Lemma 4.1
Fix $x^{\prime}\in\Omega$, with $\Omega=[0,1]^{d}$. Let $x=(x_{1},\dots,x_{d})\in\Omega$. It is obvious if any $x_{1},\dots,x_{d}$ is 0 or $1$, then $k(x,x^{\prime})=0$. To show $k\in H^{1}(\Omega)$, define the partial sum
$k^{S}(x,x^{\prime})=2^{d}\sum_{|\alpha|=1}^{S}\frac{\sin(\alpha\pi x)\sin(\alpha\pi x^{\prime})}{\pi^{2}|\alpha|^{2}},$
Because the convergence is absolute and uniform, $\exists M>0$ such that for any $x\in\Omega$,
$2^{d}\sum_{|\alpha|\in\mathbb{N}}\frac{|\sin(\alpha\pi x)\sin(\alpha\pi x^{\prime})|}{\pi^{2}|\alpha|^{2}}\leq M.$
Now for $x,y$ in $\Omega$,
$|k^{S}(x,x^{\prime})-k^{S}(y,x^{\prime})|$ $=\left|2^{d}\sum_{|\alpha|=1}^{S}\frac{\left[\sin(\alpha\pi x)-\sin(\alpha\pi y)\right]\sin(\alpha\pi x^{\prime})}{\pi^{2}|\alpha|^{2}}\right|$
$\leq 2^{d}\sum_{|\alpha|=1}^{S}\left|\frac{\left[\sin(\alpha\pi x)-\sin(\alpha\pi y)\right]\sin(\alpha\pi x^{\prime})}{\pi^{2}|\alpha|^{2}}\right|$
$=2^{d}\sum_{|\alpha|=1}^{S}\frac{\left|\sin(\alpha\pi x)-\sin(\alpha\pi y)\right||\sin(\alpha\pi x^{\prime})|}{\pi^{2}|\alpha|^{2}}$
$\leq 2^{d+1}\sum_{|\alpha|=1}^{S}\frac{|\sin(\alpha\pi x^{\prime})|}{\pi^{2}|\alpha|}\|x-y\|\leq 2M\|x-y\|.$
As the Lipschitz constant of $k^{S}$ is bounded for any $S$, and $\lim_{S\to\infty}k^{S}=k$ uniformly, $k$ is also Lipschitz continuous. Hence, $k$ is weakly differentiable. We have shown that $k\in H_{k}$, satisfying property (i) of Definition 2.1.
$\langle u,k(\cdot,x^{\prime})\rangle_{H_{k}}=\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\langle u,\psi_{\alpha}\rangle\langle k(\cdot,x^{\prime}),\psi_{\alpha}\rangle,$
where $\psi_{\alpha}$ is any orthonormal basis. Pick the basis to be the d-dimensional Fourier sine series $\psi_{\alpha}=2^{d/2}\sin(\alpha\pi x)$. We can expand $u$ by $u=\sum_{\alpha\in\mathbb{N}^{d}}\langle u,\psi_{\alpha}\rangle\psi_{\alpha}$. Now, we have for any fixed $\alpha$
$\left\langle k(\cdot,x^{\prime}),2^{d/2}\sin(\alpha\pi\cdot)\right\rangle$ $=\int\left(\sum_{|\gamma|\in\mathbb{N}}2^{d}\lambda_{\gamma}\sin(\gamma\pi x)\sin(\gamma\pi x^{\prime})\right)\left(2^{d/2}\sin(\alpha\pi x)\right)dx$
$=\int\sum_{|\gamma|\in\mathbb{N}}\left\{2^{3d/2}\lambda_{\gamma}\sin(\gamma\pi x)\sin(\alpha\pi x)\sin(\gamma\pi x^{\prime})\right\}dx$
$=\sum_{|\gamma|\in\mathbb{N}}\left\{2^{d/2}\lambda_{\gamma}\sin(\gamma\pi x^{\prime})\int 2^{d}\sin(\gamma\pi x)\sin(\alpha\pi x)dx\right\}$
$=2^{d/2}\lambda_{\alpha}\sin(\alpha\pi x^{\prime}),$
where the last line holds as $\langle 2^{d}\sin(\gamma\pi x),\sin(\alpha\pi x)\rangle=1$ for $\alpha=\gamma$ and 0 otherwise (it is the orthonormal basis we picked).
$\langle u,k(\cdot,x^{\prime})\rangle_{H_{k}}$ $=\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\langle u,\psi_{\alpha}\rangle\langle k(\cdot,x^{\prime}),\psi_{\alpha}\rangle$
$=\sum_{|\alpha|\in\mathbb{N}}\lambda_{\alpha}^{-1}\langle u,\psi_{\alpha}\rangle 2^{d/2}\lambda_{\alpha}\sin(\alpha\pi x^{\prime})$
$=\sum_{|\alpha|\in\mathbb{N}}\langle u,\psi_{\alpha}\rangle 2^{d/2}\sin(\alpha\pi x^{\prime})$
$=\sum_{|\alpha|\in\mathbb{N}}\langle u,\psi_{\alpha}\rangle\psi_{\alpha}(x^{\prime})=u(x^{\prime}).$
This shows requirement (ii) of Definition 2.1 also holds, and $k$ is the unique reproducing kernel for $H_{k}$.
Appendix C Approximation in the Schwartz space
While Theorem 4.4 is valid only in $H^{-\tau}$ (so that the Gaussian measure is well-defined), we can provide a similar convergence condition in the setting of Lévy white noises (generalized random fields) for $d>1$, provided that we let $\Omega=\mathbb{R}^{d}$. The drawback of working with the space $H^{-\tau}$ is that one must choose a specific value of $\tau$. Thus, the stochastic process described by the prior does not inherently live on a single separable Hilbert space, and a choice must be made on the underlying ambient space. In contrast, when working with Lévy white noises, there is no need to worry about ensuring that the covariance form is trace-class. This choice is perhaps the most natural or intrinsic way to view the prior, as there is no need to pick a specific $\tau$, as the characterization as a Lévy white noise always exists. The drawback, however, is that in this viewpoint, the prior is not a Gaussian probability measure, hence there is no natural Wasserstein metric and Bayes’s theorem no longer holds. Note that GP regression is still possible in this situation.
Recall the Schwartz space of smooth, rapidly decaying functions
Let $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ denote the dual of $\mathcal{S}(\mathbb{R}^{d})$. Note that $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is known as the space of tempered distributions, due to the test function topology of $\mathcal{S}\left(\mathbb{R}^{d}\right)$. We can rely on Lévy’s continuity theorem [bierme2017generalized] in the setting of tempered distributions. This yields a parallel convergence theorem to Theorem 4.4 in $\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
Theorem C.1.
Let $d>1$, $(\psi_{i})_{i\in\mathbb{N}}$ be an orthonormal basis for $L^{2}(\mathbb{R}^{d})$, and for each $n\in\mathbb{N}$, let $\mu_{n}$ be given by eq. (4.2). Then there exists a Lévy white noise $\mu$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ such that $\mu_{n}$ converges in distribution to $\mu$ under the strong topology. Further, $\mu$ has characteristic function $\phi_{\mu}(u)=\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}$, $u\in\mathcal{S}(\mathbb{R}^{d})$, hence $\mu$ can be regarded as a Gaussian random field over $\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
Remark C.1.
The space $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is inconveniently large in a machine learning context. For instance, the Dirac delta distribution belongs to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, and we cannot make sense of generating such a sample. However, there are often smaller function spaces with full measure under a Lévy white noise which are easier to characterize. A general methodology for identifying a Besov space where this property holds is provided in [aziznejad2020wavelet]. In this case, the distribution is simply the $d$-dimensional Brownian sheet, conditioned to be zero on the boundaries. One can show this process is continuous on the unit cube [adler2010geometry].
@@ Appendix C Approximation in the Schwartz space
Definition C.1 (Random field, Lévy white noise).
Let $(\Omega,\mathcal{F},\mu)$ be a probability space, and $U\subseteq\mathbb{R}^{d}$ an open set.
  1. 1.
    A random field $X$ on $U$ is a measurable mapping $X:U\times\Omega\to\mathbb{R}^{n}$ such that for any $x\in U$, $X(x;\cdot)$ is a real-valued random vector.
  2. 2.
    A Lévy white noise $X$ (generalized random field) is a measurable mapping $X:\mathcal{S}^{\prime}(\mathbb{R}^{d})\times\Omega\to\mathbb{R}^{n}$. That is, $X(u)$ is a real-valued random vector $\forall u\in\mathcal{S}(\mathbb{R}^{d})$.
The characteristic function of a Lévy white noise is defined as the Fourier transform of $X$, i.e.,
$\phi_{X}(u)=\mathbb{E}[\exp\{iX(u)\}]=\int_{\mathcal{S}^{\prime}(\mathbb{R}^{d})}\exp\{iL(u)\}d\mu_{X}(L),\quad u\in\mathcal{S}(\mathbb{R}^{d}).$
The main idea of the proof is to apply the following version of Lévy’s continuity theorem in the setting of tempered distributions.
Theorem C.2 (Lévy’s continuity theorem [bierme2017generalized]).
Let $(X_{n})_{n\in\mathbb{N}}$ be a collection of Lévy white noises, each with characteristic function $\phi_{X_{n}}$. Suppose that there exists a function $\phi:\mathcal{S}(\mathbb{R}^{d})\to\mathbb{C}$, which is continuous at 0, such that $\phi_{X_{n}}\to\phi$ pointwise. Then there exists a Lévy white noise $X$, with characteristic function $\phi$, such that $X_{n}\overset{d}{\to}X$ under the strong topology.
To apply Theorem C.2, we must show that the finite-dimensional measure given by eq. (4.2) admits a Lévy white noise representation. For this, we can rely on the Minlos-Bochner theorem.
Theorem C.3 (Minlos-Bochner theorem [bierme2017generalized]).
Let $\phi:\mathcal{S}(\mathbb{R}^{d})\to\mathbb{C}$ with $\phi(0)=1$ be positive-definite and continuous at 0. Then there exists a Lévy white noise $X$ defined on some probability space $(\Omega,\mathcal{F},\mu)$ such that $\phi$ is the characteristic function of $X$.
Remark C.2.
We clarify the precise meaning of the convergence appearing in Theorem C.2. We will use $L$ to denote elements of $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. The strong topology $\tau$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is generated by the collection of semi-norms
$q_{B}(L)=\sup_{u\in B}|L(u)|,\quad B\subset\mathcal{S}(\mathbb{R}^{d}),$
where $B$ is bounded. Now, let $(X_{n})_{n\in\mathbb{N}}$ and $X$ be Lévy white noises with measures $(\mu_{n})_{n\in\mathbb{N}}$ and $\mu$. We say that $X_{n}$ converges to $X$ in distribution under the strong topology, denoted by $X_{n}\overset{d}{\to}X$, if
$\lim_{n\to\infty}\int_{\mathcal{S}^{\prime}(\mathbb{R}^{d})}F(L)d\mu_{X_{n}}(L)=\int_{\mathcal{S}^{\prime}(\mathbb{R}^{d})}F(L)d\mu(L),\quad\forall F\in\mathcal{C}_{b}(\mathcal{S}^{\prime}(\mathbb{R}^{d}),\tau),$
with $\mathcal{C}_{b}(\mathcal{S}^{\prime}(\mathbb{R}^{d}),\tau)$ being the space of bounded continuous forms on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ under $\tau$.
We can now prove the following.
Lemma C.1.
For each $\mu_{n}$ as given by eq. (4.2), the associated random field $X_{n}$ admits a version which is a Lévy white noise.
@@
Restrict each $\mu_{n}$ to $\mathcal{S}(\mathbb{R}^{d})$, which can be done as $\mathcal{S}(\mathbb{R}^{d})\subset L^{2}(\mathbb{R}^{d})$. Then, each $\mu_{n}$ is a Gaussian measure associated with a Gaussian random field $X_{n}$ on $\mathcal{S}(\mathbb{R}^{d})$. Since $X_{n}$ is Gaussian, it has characteristic function
$\phi_{X_{n}}(u)=\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}_{n}}\hat{u}\rangle\right\},\quad u\in\mathcal{S}(\mathbb{R}^{d}),$
where $\hat{u}=\sum_{j=1}^{n}\langle u,\psi_{j}\rangle\psi_{j}$ given an orthonormal basis $(\psi_{j})_{j=1}^{n}\subset\mathcal{F}_{n}$. Observe that $\phi_{X_{n}}(0)=1$ and note that $\phi_{X_{n}}$ is positive-definite as any characteristic function is positive-definite. We also have $\langle\cdot,\Sigma_{\mathcal{F}_{n}}\cdot\rangle=\|\cdot\|_{\Sigma_{\mathcal{F}_{n}}}^{2}$, and recalling that any norm is continuous, we identify that $\phi_{X_{n}}$ is a composition of continuous functions. Hence $\phi_{X_{n}}$ is continuous at 0, and application of the Minlos-Bochner theorem completes the proof. ∎
Finally, we will look to apply Theorem C.2. By Lemma C.1, we may regard each $\mu_{n}$ in the sequence as a Lévy white noise. As they are Gaussian, each has characteristic function
$\phi_{X_{n}}(u)=\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}_{n}}\hat{u}\rangle\right\},\quad u\in\mathcal{S}(\mathbb{R}^{d}).$
Let $\phi_{\mu}(u)=\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}$, which is continuous at 0. Then in the limit, for any $u\in\mathcal{S}(\mathbb{R}^{d})$, we have
@@
By Lévy’s continuity theorem, there exists a Lévy white noise $\mu$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ with characteristic function $\phi_{\mu}$ such that $\mu_{n}\overset{\mathrm{d}}{\rightarrow}\mu$. Finally, $\phi_{\mu}$ is the form of a characteristic function of a Gaussian random field on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, which completes the proof of Theorem C.1.